OpenStudy (anonymous):

Sam has a total of 58 DVD's and CD's. If the number of CD's is two more than three times the number of DVD's, how many CD's does he have?

OpenStudy (anonymous):

dvd + cd = 58 cd + 2 = 3 * dvd subtract these two equastions dvd - 2 = 58 - 3 * dvd dvd = 60 - 3 * dvd 4 * dvd = 60 dvd = 15 so cd = 43

OpenStudy (anonymous):

it has to either 42 or 44 not 43

OpenStudy (anonymous):

it should be 43 cds and 15 dvds. both of those satisfy the two equations I wrote on the top

OpenStudy (anonymous):

idk what i got was 44

OpenStudy (anonymous):

Let D be the number of DVDs. Then C, the number of CDs is C = 3D + 2 The total is 58, so C + D = 58 3D + 2 + D = 58 4D = 56 D = 14 C = 3(14) + 2 = 42 + 2 = 44 So he has 14 DVDs and 44 CDs